A random variable x is normally with mean 10 and variance 6.25 what is the value of p(x>12.25)


1
Expert's answer
2022-05-03T16:21:18-0400
"P(X>12.25)=1-P(Z\le\dfrac{12.25-10}{\sqrt{6.25}})"

"=1-P(Z\le0.9)\approx0.184060"